SciOp

EUROPEAN SOUTHERN OBSERVATORY

SciOp

La Silla - Science Operation Department

FEROS on the MPG/ESO-2.20m
Version: 1.1 2002-11-25 J.D.Pritchard

Highlights of the Commissioning of FEROS at the ESO/MPI 2.2m telescope

Commissioning I -- September 21 - October 17, 1998
Commissioning II -- November 11 - 30, 1998
Commissioning III -- October 22/23 - 25/26 2002
Commissioning IV -- November 13/14 - 14/15 2002
Commissioning V -- March 08/09 - 10/11 2003

Commissioning III -- October 22/23 - 25/26 2002

These are the main actions carried out during Commissioning-III. Weather and technical problems did not help, so only 3 out of the 7 Commissioning and 3 Technical nights could be use for scientifically useful observations (plus IS had to spend 1 day in bed with fever...).

The general idea is to follow the guidelines of the 1998 Commissioning . Please find below the list of targets.

Target list

Summary

FEROS is functional, but with limited capacity. The efficiency is down by a factor or ~3 (compared to to the benchmark of 15% set at the ESO-1.52m).

Log

Commissioning 1998 vs. Commissioning-III 2002

De-blazing

Tau Ceti
Tau Ceti new
Above we show a DRS-reduced, 180 sec spectrum of tau Ceti, taken in OC mode. In particular, orders #17 and #18 around Hbeta are plotted (see Table 1 for the wavelength coverage vs. order number). The S/N of the two spectra is comparable, though the DRS did a slightly better job in deblazing the original spectrum.

Spectral coverage

SMC R50 was used to explore the spectral coverage of FEROS, as it was done originally. Below we plot the spectrum of that star, in four wavelength ranges. For each panel, the left graph comes from the original commissioning, the right graph is that obtained in October 2002. The original observation is a 5400 sec exposure, our observation is an average of 3x1800 sec exposures. Evidently, our spectrum has lower S/N, however the seeing was oscillating around 1.5" fwhm, and we now have a 1.8" fiber diameter on the sky.

Region of Hdelta
SMC R50
SM50 Hdelta


Region of HeI 5876 (in the right panel, the spike at the center of the HI line is a cosmic)
SM50
Helium 5876


Region of Halpha
SMC R50
SMC R50 Halpha


Region of Paschen 13
SMC R50
SMC R50 Paschen 13

Maximum S/N

A 15 sec spectrum of beta Orionis (V=0.12) gives S/N=120, it was S/N=700 during CommI. After the commissioning was completed the fibre head was examined and it was found that the alignment of the fibre apertures, ball lenses and fibres was far from optimal. So at the moment it is beleived that this is the main cause of the low efficiency, and of the low relative fibre efficiency since the alignmnet for the sky fibre
is noticably worse than for the object fibre.

Efficiency

The ca. 15% peak efficiency measured at the 1.52m was never reached, due to the poor quality of the night, and the non-optimal alignment of the fiber-entrance--ball microlens--fiber pupil (see above). A maximum around 8% was measured in the morning of Oct 26, and that is plotted below. The standard star HR1544 was used. The seeing was down to around 1" fwhm, but the airmass was 1.4. Moreover, a wrong effective area was used in the computation, so the real peak efficiency is around 6.6%. See this email by AK, and my reply.

HR 1544

Clocks

The FITS header of a few images was checked. On Oct 26, the coordinate info (both apparent and from catalog) were correct, and so were the airmass, ZD, and hour angle. The UTs were a more mixed situation, as follows:

Starting an exposure at UT 6:22:36

in the header it was:

TM_START = 6:22:17
TCSUTC = 6:22:34

so TM_START was late by 19 seconds, TCSUTC was about right.

Apropos the workstations, I found

UT(CERME) - UT(bias) = 4 sec
UT( off) - UT(bias) = 2 sec

Tables

Table 1: FEROS wavelength coverage vs. order
Order #   wavelength  [Å]

   
0001 -- 3528-3661
0002 -- 3584-3720
0003 -- 3643-3781
0004 -- 3703-3844
0005 -- 3766-3909
0006 -- 3831-3976
0007 -- 3898-4046
0008 -- 3967-4118
0009 -- 4039-4193
0010 -- 4114-4270
0011 -- 4191-4351
0012 -- 4272-4434
0013 -- 4355-4521
0014 -- 4442-4611
0015 -- 4533-4705
0016 -- 4627-4803
0017 -- 4726-4905
0018 -- 4828-5012
0019 -- 4935-5123
0020 -- 5047-5239
0021 -- 5165-5361
0022 -- 5288-5488
0023 -- 5416-5622
0024 -- 5552-5762
0025 -- 5694-5910
0026 -- 5844-6065
0027 -- 6002-6229
0028 -- 6168-6402
0029 -- 6344-6585
0030 -- 6531-6778
0031 -- 6729-6984
0032 -- 6939-7202
0033 -- 7163-7434
0034 -- 7401-7682
0035 -- 7657-7946
0036 -- 7930-8230
0037 -- 8224-8535
0038 -- 8540-8863
0039 -- 8881-9217

Commissioning IV -- November 13/14 - 14/15 2002 + November 15/16 - 24/25 half nights of reserved time...

Summary

FEROS is now fully operational, though the efficiency is still down compared to the ESO-1.52m by a factor of ~1.5.

Log

Commissioning 1998 vs. Commission-IV 2002

Efficiency

The ca. 15% peak efficiency measured at the 1.52m has still not been achieved. It is beleived that the remaining misalignment between the Entance Aperture in the fibre head and the MicroLens+Fibre assembly is resposible. The maximum efficiency measured thus far is shown below, with a peak efficiency of ~11.5% which is to be compared with 6.6% from Commissioning-III, i.e. an improvement by a factor of ~2 resulting from the improved alignment of the MicroLens+Fibre assembly achieved during the time between Comm-III and Comm-IV. Raw Image FEROS.2002-11-16T01:35:34.000.fits, object HR9087, exposure time 60.0sec.

HR 1544

Relative Fibre Efficiency

The Relative Fibre Efficiency computed with twilight sky flat field, and @@ fiberratio. Fiber One (object) gets ~1.2-1.5 times more flux than fiber Two (sky). Raw Image FEROS.2002-11-24T22:29:47.000.fits, object Twilight sky, exposure time 60.0sec.

FibreRatio 2002-11-24

Maximum S/N

A 5 sec spectrum of beta Orionis (V=0.12) gives S/N=200 implying 15sec would give S/N~600 however 15sec resulted in some saturation. It was S/N=700 during Comm-I and 120 during Comm-III. Raw Image FEROS.2002-11-15T08:11:57.000.fits, object ß Orionis, exposure time 5.0sec.

Resolving Power

The nominal Spectral Resolution (R) is 48000. This corresponds to a FWHM of the lines of ~2.2 pixels. The following pair of plots shows that the nomianl FWHM is achieved and moreover the spectrograph is better focused than it was at the ESO-1.52m with a flatter distribution of focus values over the dispersion direction of each order AND over the full spectral range of the spectrograph.
ThAr Focus 2002-04-01
ThAr Focus 2002-11-25


De-blazing

Tau Ceti
Tau Ceti new
Above we show a DRS-reduced, 180 sec spectrum of tau Ceti, taken in OC mode. In particular, orders #17 and #18 around Hbeta are plotted (see Table 1 for the wavelength coverage vs. order number). The S/N of the two spectra is comparable, though the DRS did a slightly better job in deblazing the original spectrum.

Spectral coverage

SMC R50 was used to explore the spectral coverage of FEROS, as it was done originally. Below we plot the spectrum of that star, in four wavelength ranges. For each panel, the left graph comes from the original commissioning, the right graph is that obtained on November 17 2002. Both the original and the 17/11/2002 observations are 5400sec exposures. Raw Image FEROS.2002-11-18T02:16:35.000.fits, object SMC R50, exposure time 5400.0sec.

Region of Hdelta
SMC R50
SM50 Hdelta


Region of HeI 5876 (in the right panel, the spike at the center of the HI line is a cosmic)
SM50
Helium 5876


Region of Halpha
SMC R50
SMC R50 Halpha


Region of Paschen 13
SMC R50
SMC R50 Paschen 13
 

Clocks

The clocks on the BIAS PCs are supposed to be synchronised with UTC using the NTP daemon. However for unknown reasons this service has not been performing correctly. Instead as an interrim solution we synchronise the BIAS PC clock to the 2.20m time server w2p2tcs using ntpdate every ten minutes (via a cronjob). The result is that the clock then drifts by only approximately 0.033±0.002secs by the end of any given 10min period.

Commissioning V -- March 08/09 - 10/11 2003 + Reserved nights March 15/16 - 17/18 2003

Summary

Efficiency slightly improved compared to prior to intervention but still must be restored to nominal 15-20%.
This (mini) commissioning was required by the following changes carried out during the March-2003 FibreHead Upgrade Mission:
  1. FibreHead Upgrade
  2. Shorter Calibration Fibres (were 20m, now 15m): Will possibly improve throughput at shorter wavelengths.
  3. FCU Fibre input unit lenses moved closer together: Should improve calibration lamp intensities as the lenses will now be closer to the center of the beam comming from the FCU.
  4. NDFW in convergent beam: Should eliminate fringing
  5. NDFW moved to new motor assembly with higher resolution encoder.
  6. FCU Fibre input unit shutter now on former NDFW motor assembly.
  7. Small Format TCCD coupled with new FFH-TCCD Optics

Log

Efficiency

The maximum efficiency measured thus far is shown below, with a peak efficiency of ~13% which is to be compared with 6.6% from Commissioning-III and 11.5% for Commissioning-IV. Raw Image FEROS.2003-03-11T08:52:35.000.fits, object HR4963, exposure time 180.0sec.

HR 4963

Relative Fibre Efficiency

The Relative Fibre Efficiency computed with twilight sky flat field, and @@ fiberratio. The throughput of fibres is almost exactly equal now. Raw Image FEROS.2003-03-16T22:04:39.000.fits, object Twilight sky, exposure time 120.0sec.

FibreRatio 2003-03-16

Resolving Power

Nominal Spectral Resolution (R) of 48000 is maintained.
Instrument Focus

Clocks

The clocks on the BIAS PCs now accurately synchronised with UTC on w2p2tcs using NTP.


Last revision: J.D.Pritchard 2003-03-17
Comments
[ESO |LaSilla |SciOp ]